OpenStudy (anonymous):

if x, y, and z are positive numbers and 2x=3y=4z, then the value of x+y+z is how many times the value of x? and why?

OpenStudy (greencat):

\[x=z\] \[z=\frac{1x}{2}\] \[y=\frac{2x}{3}\]

OpenStudy (greencat):

\[x+y+z=x+\frac{ 1x }{ 2 }+\frac{ 2x }{ 3 }\]

OpenStudy (anonymous):

thank you so much =)

OpenStudy (greencat):

Don't forget to add the fractions.

OpenStudy (greencat):

Here is a link to help you get around in openstudy (just in case) Don't forget to medal the guy who made tutorial. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

OpenStudy (greencat):

Do you need help getting the final answer?

OpenStudy (anonymous):

is it 13/6?

OpenStudy (greencat):

Don't forget x. :) \[\frac{13x}{6}\]

OpenStudy (anonymous):

thank you so much =)

jimthompson5910 (jim_thompson5910):

Another way to get the same answer W = x+y+z W = 12(x+y+z)/12 W = (12x+12y+12z)/12 W = (6*2x+4*3y+3*4z)/12 W = (6*2x+4*2x+3*2x)/12 W = (12x+8x+6x)/12 W = (26x)/12 W = (13/6)*x

OpenStudy (anonymous):

that is really helpful as well thank you =)

OpenStudy (anonymous):

Where did you get the 12 from?

OpenStudy (kamibug):

\(\huge\cal\color{green}{Welcome~to~Openstudy!~:)}\)